A 32−Ω resistor and a 20−Ω resistor are connected in parallel, and the combination is connected across a 240-V dc line. (a) What is the resistance of the parallel combination? (b) What is the total current through the parallel combination? (c) What is the current through each resistor?


1
Expert's answer
2021-06-09T08:44:29-0400

"R_1=32\Omega"


"R_2=20\Omega"

V=240V

Req="\frac{R_1\times R_2}{R_1+R_2}"

"R_{eq}=\frac{32\times20}{32+20}"

"R_{eq}=12.30\Omega"

Current

"I=\frac{V}{R_{net}}=\frac{240}{12.39}"

"I=19.51Amp"

"R_1" Flow current

"I_1=\frac{R_2}{R_1+R_2}\times I"

"I_1=\frac {20}{52}\times19.51" =7.50A

"I_2=\frac{R_1}{R_1+R_2}\times I"

"I_2=\frac {32}{52}\times19.51" =12.00A